# Dynamic watermarks in images generated by diffusion models

### v1

@@ 4 Watermark Extraction Results
To evaluate the effectiveness of our dual watermarking approach, we present the extracted watermarks from both branches: the model watermarking branch and the image watermarking branch.
4.1 Model Watermarking Branch
As shown in Fig. 4, our method effectively preserves the watermark’s visibility and structure after extraction. The central logo undergoes minor deformation due to the diffusion process’s inherent noise and transformations, which does not impact the QR code’s scalability or embedded information.
Figure 4: Extracted QR code watermark from the model watermark embedding branch. IP address and time are just hypothetical examples.
To evaluate the effectiveness of the extracted watermark, we analyze the Peak Signal-to-Noise Ratio (PSNR) (Korhonen & You, 2012) and SSIM between the original and extracted watermarks, comparing them with several mainstream models. Higher PSNR and SSIM values indicate better preservation of visual quality. As shown in Table 1, while our model does not achieve the highest PSNR and SSIM scores, it performs above average across these metrics, effectively balancing visual quality and watermark robustness.
Table 1: Comparison Based on PSNR and SSIM
Metric DCT-DWT SSL Watermark FNNS LDM DiffEdit HiDDeN Ours
PSNR 39.5 31.1 32.1 34.0 31.2 32.0 35.67
SSIM 0.97 0.86 0.90 0.94 0.92 0.88 0.89
4.2 Image Watermark Embedding Branch
As shown in Figure 5, our custom-designed ”Diffusion Model” and ”Watermark” serve as the core information. Each extracted watermark retains this text, while dynamic transformations modify its appearance. These transformations enhance robustness and resilience, yet the classifier can still accurately verify the watermark’s origin.
Figure 5: Extracted watermarks from the image watermark embedding branch. Each watermark has been dynamically transformed in shape and color.
4.3 Impact of Watermark Quantity on Extraction Accuracy
Our implementation supports up to six dynamic watermarks, maintaining robust extraction accuracy within this limit. To quantify this, we tested configurations with 2 to 6 watermarks. The extraction accuracy for each is reported in Table 2.
Table 2: Extraction Accuracy with Increasing Number of Watermarks
Number of Watermarks 2 3 4 5 6
Extraction Accuracy (%) 100.0 100.0 100.0 98.3 86.1
4.4 Validation Based on Statistical Analysis
Table 3: Image statistics comparison results. ”Difference” refers to the percentage difference in data.Each Image statistics references:(Tsai et al., 2008),(De & Masilamani, 2013),(de Ridder, 1996),(Sebastian V et al., 2012),(Sebastian V et al., 2012),(Tahmid & Hossain, 2017),(He et al., 2005),(Frank et al., 2020),(Chan et al., 2018),(Freitas et al., 2018)(Chen et al., 2024)
GLCM GLCM Canny Variance Blur Mean Edge Entropy Sharpness Saturation Texture Image
Contrast Energy Edge Measure spectrum Histogram Strength Score Realism
Watermarked 354.80 0.19 46.41 1124.47 85.73 6.00 6.31 8836.29 63.04 5.92 1.55
Clean 371.23 0.19 48.34 1209.43 88.18 7.00 6.23 8903.23 67.00 5.65 1.42
Difference 4.63% 0.00% 4.16% 7.56% 2.86% 16.67% 1.27% 0.76% 6.28% 4.56% 8.39%
We use image statistics based on human perception as evaluation metrics. Based on the sensitivity coefficient $s$ proposed by  (Chen et al., 2024), we selected 11 image statistics measures with sensitivity values above 0.3, indicating a high sensitivity to changes in texture, edges, and frequency. Sensitivity ranges from 0 to 1, with higher values indicating greater sensitivity.
As shown in Table 3, the average change between watermarked and clean images across these metrics is approximately 5%. The GLCM Energy remains unchanged (0.00%), indicating that watermark embedding does not significantly affect image energy, which reflects texture uniformity and homogeneity. The Edge Histogram shows the largest difference (16.67%), highlighting its sensitivity to structural changes introduced by watermarking. This suggests that watermarking impacts edge distribution and frequency more than other visual characteristics, making the Edge Histogram an effective measure for detecting structural alterations.
4.5 Watermark Attack Performance
We assessed the robustness of our watermarking method against common attacks, considering only those that preserve the visual quality of the generated image. For instance, while severe cropping (over 70%) reduces watermark recognition accuracy, it introduces visible tampering, making the attack perceptible. Thus, we focus on common attacks such as rotation, blurring, texture reduction, and image compression.
@@ 4.5 Watermark Attack Performance
As shown in Table 4, in the absence of attacks, ”Watermark Presence” accuracy reached 100.00%, indicating perfect detection in undisturbed images. Texture reduction caused the largest drop (4.3%), likely due to its impact on watermark details, weakening detectability. Image compression, particularly lossy compression, also reduced accuracy to 98.50%. However, flip and crop attacks had minimal impact, maintaining accuracy at 100.00% and 99.10%, respectively.
For the ”Watermark Classification” task, accuracy for unaltered images was high (97.00%). However, attacks slightly reduced classification accuracy. Blurring and texture reduction had the greatest impact, lowering accuracy to 93.97% and 93.94%, respectively, suggesting these attacks degrade visual features, making classification more challenging. Image compression and rotation also affected classification but to a lesser extent.
@@ 5 Conclusion
In this paper, we propose a dual approach for embedding fixed QR-codes within the diffusion process and dynamic watermarks in generated images. This integration enhances intellectual property protection and traceability in generated content. The dynamic watermark undergoes feature and pixel space transformations, increasing resistance to attacks while preserving image quality. It remains intact even under rotation, blurring, and compression. Statistical validation shows dual watermarking method has minimal impact on image quality, with only a 5% variation in key metrics such as edge and texture attributes, confirming the method’s invisibility and robustness. Our work addresses critical challenges in the ethical use of AI-generated content, providing a scalable and effective mechanism for ownership verification and misuse prevention. Future research could explore extending this framework to other generative models and applications, further advancing the field of digital content security.

### v2

@@ 4 Watermark Extraction Results
Table 1: Image statistics comparison results. ”Difference” refers to the percentage difference in data.
. GLCM GLCM Canny Variance Blur Mean Edge Entropy Sharpness Saturation Texture Image Contrast Energy Edge Measure spectrum Histogram Strength Score Realism Watermarked 354.80 0.19 46.41 1124.47 85.73 6.00 6.31 8836.29 63.04 5.92 1.55 Clean 371.23 0.19 48.34 1209.43 88.18 7.00 6.23 8903.23 67.00 5.65 1.42 Difference 4.63% 0.00% 4.16% 7.56% 2.86% 16.67% 1.27% 0.76% 6.28% 4.56% 8.39%
@@
In row 1 of Fig. 3, our method effectively preserves the watermark’s visibility and structure after extraction. The central logo undergoes minor deformation due to the diffusion process’s inherent noise and transformations, which does not impact the QR code’s scalability or embedded information. In row 2, our custom-designed ”Diffusion Model” and ”Watermark” serve as the core information. Each extracted watermark retains this text, while dynamic transformations modify its appearance. These transformations enhance robustness and resilience, yet the classifier can still accurately verify the watermark’s origin.
Figure 3: Raw 1 shows the extracted QR code watermark from the model watermark embedding branch. Raw 2 shows extracted watermarks from the image watermark embedding branch
We use image statistics based on human perception as evaluation metrics. Based on the sensitivity coefficient proposed by  [7], we selected 11 image statistics measures with sensitivity values above 0.3, indicating a high sensitivity to changes in texture, edges, and frequency. Sensitivity ranges from 0 to 1, with higher values indicating greater sensitivity. As shown in Table 1, the average change between watermarked and clean images across these metrics is minimal, approximately 5%.
As shown in Table 2, texture reduction caused the largest drop (4.3%), likely due to its impact on watermark details, weakening detectability. Image compression, particularly lossy compression, also reduced accuracy to 98.50%. However, flip and crop attacks had minimal impact, maintaining accuracy at 100.00% and 99.10%, respectively. For the ”Watermark Classification” task, attacks slightly reduced classification accuracy. Blurring and texture reduction had the greatest impact, lowering accuracy to 93.97% and 93.94%, respectively, suggesting these attacks degrade visual features, making classification more challenging.
Table 3: Comparison of Image Generation Quality using IS and FID
Method IS $\uparrow$ FID $\downarrow$
Tree-Ring [27] 3.80 3298.22
Chen et al. [8] 4.43 2789.98
Xin et al. [19] 4.14 3477.38
Clean images 5.01 2157.45
Ours 4.61 2687.34
To evaluate the effectiveness of our watermark embedding method in image generation, we compare our approach with three existing watermark embedding techniques using two widely used metrics: Inception Score (IS) [25] and Fréchet Inception Distance (FID) [16]. The IS measures the diversity and realism of the generated images, where higher scores indicate better quality. The FID assesses the distribution similarity between generated and real images, with lower scores indicating better quality. The comparison results are presented in Table 3. Our method outperforms the other approaches, achieving a higher IS and a lower FID, demonstrating its superior performance in maintaining image quality while embedding watermarks.
@@ 5 Conclusion
In this paper, we propose a dual approach for embedding fixed QR-codes within the diffusion process and dynamic watermarks in generated images. This integration enhances intellectual property protection and traceability in generated content. The dynamic watermark undergoes feature and pixel space transformations, increasing resistance to attacks while preserving image quality. It remains intact even under rotation, blurring, and compression. Statistical validation shows dual watermarking method has minimal impact on image quality, with only a 5% variation in key metrics such as edge and texture attributes, confirming the method’s invisibility and robustness. Our work addresses critical challenges in the ethical use of AI-generated content, providing a scalable and effective mechanism for ownership verification and misuse prevention. Future research could explore extending this framework to other generative models and applications, further advancing the field of digital security.
6 Acknowledgement
This research was partially supported by National Intelligence and Security Discovery Research Grants (project NS220100007), funded by the Department of Defence, Australia